OpenStudy (anonymous):

The point (-1,5) is half-way between (a,-9a) and (a+3,4a). What is the value of "a"?

OpenStudy (lgbasallote):

hint: \[\huge \frac{a + (a+3)}{2} = 1\] \[\huge \frac{-9a + 4a}2 = 5\] does that help?

OpenStudy (lgbasallote):

solve for a

OpenStudy (anonymous):

for both?

OpenStudy (lgbasallote):

wait first one should be -1 not 1

OpenStudy (lgbasallote):

you can do both...but they'll just give you the same values for a

OpenStudy (lgbasallote):

pick whichever equation is easier for you then solve for a

OpenStudy (anonymous):

for the 1st i got a=-1/2 and the 2nd i got -2/3

OpenStudy (lgbasallote):

redo your solution...i think you did something wrong

OpenStudy (lgbasallote):

remember...first equation is = -1 not =1

OpenStudy (anonymous):

-5/2

OpenStudy (lgbasallote):

\[\huge \frac{a + (a+3)}{2} = -1\] cross multiply \[\huge a + a + 3 = -2\] is this what you did?

OpenStudy (lgbasallote):

something smells fishy here

OpenStudy (amistre64):

(a,-9a) ( -1 , 5 ) (a+3,4a) -a +9a -a +9a -a +9a ------------------------ (0,0) (-1-a,5+9a) (3,13a)\[-1-a=\frac{3}{2}\]

OpenStudy (anonymous):

i'm still confused

OpenStudy (amistre64):

if I lay a pencil on the desk in front of me, it has a middle point; if i pick up the pencil and move it to the table in the kitchen, its middle point is still the same relative to the pencil

OpenStudy (amistre64):

my method, i moved the pencil to see it better

OpenStudy (amistre64):

-1-a is still half way between 0 and 3

OpenStudy (anonymous):

ok